# Dark Kinetic Heating of Exoplanets and Brown Dwarfs

### v1

@@ Abstract
Dark matter heating of low-escape velocity objects has been considered, but only with dark matter annihilation heat rather than kinetic heating. One class of ideal, low-escape velocity dark matter detectors where dark matter annihilation heat has been considered includes exoplanets and brown dwarfs Leane and Smirnov (2021). These objects are ideal because they have lower core temperatures relative to $e.g.$ nuclear-burning stars, leading to retention of lighter dark matter, as well as lower heat backgrounds. Dark matter heating in exoplanets and brown dwarfs is also likely much easier to detect than neutron stars, due to their substantially larger surface areas and therefore larger potential luminosities. Exoplanets and brown dwarfs can therefore be detected far into the Galactic center, providing a potential tracer of the unknown dark matter density distribution in the Galaxy Leane and Smirnov (2021). There are a large number of candidates detected or soon to be detected, with the onslaught of new high-power infrared and optical telescopes such as JWST, Roman, and Rubin, making this search promising Leane and Smirnov (2021); Perryman et al. (2014); Green et al. (2012); Johnson et al. (2020).
@@
I Toy Model Setup
@@
where $m_{\chi}$ is the dark matter mass. It is also of course possible to consider a Standard Model coupling, as was considered previously for neutron star kinetic heating Gresham et al. (2023). However, experimental constraints on the Standard Model coupling typically lead to weaker signals for the same value of coupling $g_{\chi}$. On the other hand, $g_{\chi}$ is very weakly constrained, with the strongest constraints claimed by observations of the Bullet Cluster and halo shapes, which presently constrain $g_{\chi}\lesssim 10^{-4}\,(m_{\chi}/{\rm MeV})^{3/4}$ Knapen et al. (2017). We therefore focus on long-range dark matter self-interactions. Lastly, higher-order interactions of $\phi$ are in principle allowed as well, but we will consider scenarios where such terms are negligible. We discuss extensions to a non-zero quartic interaction in Appendix A.
@@
Once sufficient dark matter is captured by a celestial object, there are three main effects that the potential in Eq. (2) has on the heating of celestial bodies which cause the calculation to differ from the standard short-range scenario. First, there will be an increased focusing effect that causes more dark matter to pass through the object than would from purely gravitational interactions. Second, the energy that the dark matter has at the surface of the object will be greater. Finally, the dark matter can self-scatter via the long range interaction. In the presence of an additional sufficiently strong contact interaction between dark matter and the Standard Model, however, we can simplify the calculation by ignoring the third effect. For simplicity, we will work with this condition and calculate conservative sensitivities assuming the presence of this additional contact interaction. This is not necessarily required, but as the goal of our work is to demonstrate for the first time that detectable dark kinetic heating can occur in low-escape velocity objects, we choose this simple case as a starting point. We will investigate a more general form of capture without contact interactions in future work.
@@ II.2.1 Case I: Low Dark Matter Abundance, Low Strength Field
We first analyze the case of a single centrifugal barrier, when the dark potential $\Phi$ is sufficiently weak, and the force range $\lambda$ is sufficiently long, such that the $1/r$ factor in the dark potential $\Phi(r)$ limits capture. Because we are interested in a heating signal, we require that no centrifugal barrier be present at or beyond the surface of the celestial object, as otherwise this would prevent the dark matter from depositing its kinetic energy. This defines a maximum angular momentum, and therefore a maximum impact parameter $b_{\rm max}^{\rm dark}$, for which the point of closest approach matches the surface of the object ($i.e.$ Eq. (7) vanishes at the surface of the celestial object). Dark matter particles with angular momentum $b\leq b_{\rm max}^{\rm dark}$ thus contribute to heating, while those with $b>b_{\rm max}^{\rm dark}$ are unable to reach the object and deposit their energy. The full calculation is shown in Appendix B, however we predominantly work in the limit that $|\Phi|\ll m_{\chi}$. Under this approximation, we find that
@@ II.2.2 Case II: Sufficiently High Dark Matter Abundance, High Strength Field
As the number of dark matter particles inside the celestial body and therefore the dark potential grows over time, the maximum impact parameter also grows. While higher order terms in Eq. (11) need to be included in this regime, the fact that the maximum impact parameter $b_{\rm max}^{\rm dark}$ increases with increasing dark potential $\Phi$ can already be seen from Eq. (11). As the dark potential has a Yukawa scaling, its strength is exponentially suppressed at distances larger than the force range $\lambda$. Therefore, once a large amount of dark matter is accumulated and the impact parameter grows past the force range, there is eventually a shutoff where exponential growth no longer occurs, and the growth of the dark matter number becomes sub-linear. This effect is taken into account via a second centrifugal barrier, and is only relevant in the large dark matter number regime; we derive it in Appendix B. However, for most of our parameter space of interest, and most of the evolution time, Eq. (11) is sufficient and we focus our analytic calculations in the following sections to Case I.
@@ II.3 Dark Matter Energy Loss
In addition to increasing the effective capture radius, the long range force will also significantly boost in-falling dark matter particles. The boost factor as measured in the rest frame of the surface of the celestial object is given by (see Appendix C)
@@ II.4 Capture Probability
is the transition cross section, which marks the transition from the single to multi scatter regime, and $N_{\rm SM}$ is the number of target particles inside the celestial body of interest.
@@ II.5 Accumulation Rate
where $v_{p}$ is the peak of the dark matter velocity distribution (see Appendix D for details). We can now readily integrate Eq. (22) by writing
@@
where we have assumed the initial condition $N_{\chi}(0)=0$. Note that this equation fully relies on using Eq. (11) for $b_{\rm max}^{\rm dark}$, and as such, only applies to Case I of Section II.2. This is an exponential accumulation phase, given the form of Eq. (27). At very late times (or very large coupling $\alpha$ relative to force length $\lambda$), eventually the exponential growth of $N_{\chi}$ is shut off; see Appendix B.
In our considered parameter space, heating occurs before an accumulated dark matter number greater than about $10^{-6}N_{\rm SM}$. Once the dark matter density becomes comparable to that of the Standard Model density, other relevant changes to the object’s behavior could occur. Furthermore, once the dark matter number is high enough, the dark potential becomes comparable to the dark matter mass, and additional physical effects may become relevant that are outside the scope of this work ($e.g.$ pair production of dark matter particles and rapid radiation of the mediator, see Appendix C). For these reasons, we analyze only the scenario where $|\Phi|\lesssim m_{\chi}$. We find observable heating before this point, and note that for stronger couplings there will certainly be some observable signal, even if different to that from an increase in temperature pointed out here.
@@ III.1 Instantaneous Heating
with $T_{\rm SM}$ the temperature expectation from Standard Model sources in the absence of dark matter heating.
@@ IV Exoplanet and Brown Dwarf Kinetic Heating
Equipped with a framework for dark kinetic heating in the presence of long-range forces, we now apply our calculation to an example class of low-escape velocity objects: exoplanets and brown dwarfs. As elucidated in the introduction, these objects are advantageous for detection due to their large surface areas, leading to larger luminosities for fixed temperatures, but they also have an extensive upcoming search program with $e.g.$ JWST, Roman, and Rubin Telescopes. Tests far into the Galactic center also may allow potential probes into the largely unknown dark matter distribution in our Milky Way Galaxy. However, they have previously only been studied in the context of dark matter annhilation heating Leane and Smirnov (2021). We will show now that the discovery potential for this search is greatly enhanced when considering broader classes of particle physics models that produce the dark kinetic heating effect. We will consider other low-escape velocity objects in Sec. V that, depending on the model parameters, may also be of complementary interest.
@@ IV.1 Sensitivity to Coupling Strength and Force Range
Figure 4: Dark kinetic heating of a 10 Gyr old, 55 Jupiter mass brown dwarf with Jupiter’s radius, at local position as a function of the force range $\lambda$, with fixed coupling strength $\alpha=3\times 10^{-27}$. The gray dashed line “bg” indicates a background temperature of 750K, the blue line shows the dark kinetic heating, and the orange line displays the total brown dwarf temperature.
Figure 4 demonstrates dark kinetic heating of a brown dwarf as a function of force range $\lambda$. Here we have again taken $m_{\chi}=1$ GeV for simplicity. We choose a fixed value of the coupling $\alpha=3\times 10^{-27}$ as we vary the force range. This value is chosen as the minimum value of $\alpha$ such that heating asypmtotes when $N_{\chi}\lesssim\mathcal{O}(10^{-6})\times N_{\rm SM}$. We emphasize that this coupling $\alpha$ is chosen for demonstrative purposes, and any $\alpha\gtrsim 3\times 10^{-27}$ would have strong heating signals as well. This plot shares features similar to Fig. 3, where the left side asymptotes to the same value in the limit of negligible dark potential, and then an exponential increase in heating occurs as we increase the force range. Unlike the case of increasing the coupling $\alpha$, however, heating eventually saturates to a set value as we increase the force range large enough. This is because after a certain range ($\lambda\sim\mathcal{O}(10^{7})$ km for the example parameters we have chosen), the long range force is already behaving as essentially infinite relative to other parameters of the theory.
@@ IV.4 Constraints from WISE and JWST Data on a Super-Jupiter Exoplanet
Intriguingly, for dark matter parameters that would give the exact amount of heating required to sustain the 250 K temperature, one could also question whether the Super-Jupiter WISE 0885-0714 is actually significantly older and represents a positive dark matter signal, for alternate dark matter parameters. Indeed, it is important to note that the mass, temperature, and age of this Super-Jupiter are not measured independently, but rather inferred based on spectral data Leggett et al. (2017). This inference includes Standard Model-only assumptions, and therefore would change under the addition of dark kinetic heating. In any case, limits can however be drawn by taking conservative choices for any inferred Standard Model properties.
@@
Figure 8: Cross section required for 99% of dark matter particles that pass through a 55 Jupiter mass brown dwarf with with the radius of Jupiter to be captured. An example fixed dark matter number of $N_{\chi}=5\times 10^{-12}N_{\rm SM}$ is used here, where $N_{\rm SM}$ is the number of Standard Model nucleons in the object. A fixed force range $\lambda$ equal to 100 times the brown dwarf’s radius is also chosen. The dashed blue line represents dark interaction strengths $\alpha=0$, displaying the required cross section in the absence of a long range force, and the solid lines represent different values of $\alpha$. The values of $N_{\chi}$ and $\alpha$ are arbitrarily chosen to demonstrate the effect on $\sigma_{\rm max}$ of a long range force, as $N_{\chi}$ is a function of time for any chosen value of $\alpha$.
@@ IV.5 Scattering Cross Section Sensitivity
Lastly, note in Figs. 8 and 9 that we have ignored any potential evaporation effects. Evaporation occurs when the dark matter is up-scattered by the celestial body matter, such that it can overcome the object’s escape velocity and leave the object, removing any potential dark matter signal. While long-range self-interactions will mitigate evaporation once some dark matter has been accumulated, the rapid evaporation of very light dark matter might negate the possibility of any build-up at all. However, there are a wide range of scenarios where evaporation does not generally occur, even at very light dark matter masses. For example, when the field $\phi$ also couples to nucleons evaporation is greatly mitigated Acevedo et al. (2024b), and can be suppressed by multiple orders of magnitude below usual contact interaction expectations. We expand on this point in Appendix E. We therefore simply note that some regions of Fig. 9 might require more specific ingredients than the minimal toy model adopted here.

### v2

@@ Abstract
Dark matter heating of low-escape velocity objects has been considered, but only with dark matter annihilation heat rather than kinetic heating. One class of ideal, low-escape velocity dark matter detectors where dark matter annihilation heat has been considered includes exoplanets and brown dwarfs Leane and Smirnov (2021). These objects are ideal because they have lower core temperatures relative to $e.g.$ nuclear-burning stars, leading to retention of lighter dark matter, as well as lower heat backgrounds Leane and Smirnov (2021); Leane and Tong (2024). Dark matter heating in exoplanets and brown dwarfs is also likely much easier to detect than neutron stars, due to their substantially larger surface areas and therefore larger potential luminosities. Exoplanets and brown dwarfs can therefore be detected far into the Galactic center, providing a potential tracer of the unknown dark matter density distribution in the Galaxy Leane and Smirnov (2021); Benito et al. (2024). There are a large number of candidates detected or soon to be detected, with the onslaught of new high-power infrared and optical telescopes such as JWST, Roman, and Rubin, making this search promising Leane and Smirnov (2021); Perryman et al. (2014); Green et al. (2012); Johnson et al. (2020); Leane and Tong (2024); Benito et al. (2024).
@@
I Particle Physics
I.1 Toy Model Setup
@@
where $m_{\chi}$ is the dark matter mass. It is also of course possible to consider a Standard Model coupling to $\phi$, as was considered previously for neutron star kinetic heating Gresham et al. (2023). However, experimental constraints on the Standard Model coupling typically lead to weaker signals for the same value of coupling $g_{\chi}$. On the other hand, $g_{\chi}$ is very weakly constrained, with the strongest constraints claimed by observations of the Bullet Cluster and halo shapes, which presently constrain $g_{\chi}\lesssim 10^{-4}\,(m_{\chi}/{\rm MeV})^{3/4}$ Knapen et al. (2017). We therefore focus on long-range dark matter self-interactions. Lastly, higher-order interactions of $\phi$ are in principle allowed as well, but we will consider scenarios where such terms are negligible. We discuss extensions to a non-zero quartic interaction in Appendix B.
@@
I.2 Physical Effects and Motivations
Once sufficient dark matter is captured by a celestial object, the potential in Eq. (2) introduces three primary effects on heating that distinguish the calculation from the standard short-range scenario. First, the long-range force enhances gravitational focusing, causing more dark matter to intersect the object than would occur through gravity alone. Second, dark matter arrives at the surface with increased kinetic energy due to the additional potential. Third, the dark matter may undergo self-scattering via the long-range interaction. In the presence of an additional sufficiently strong contact interaction between dark matter and the Standard Model, however, we can simplify the calculation by ignoring the third effect. For simplicity, we adopt this assumption and calculate conservative sensitivities by assuming a large enough cross section to facilitate capture. While some interaction with baryons is necessary for observable heating, it need not dominate the capture process. Our goal is to establish, for the first time, that detectable dark kinetic heating can occur in objects with low escape velocities. As such, we begin with this simplified case and leave a general treatment for future work.
Another possible physical effect is the formation of multi-particle bound states. However, for the couplings and interaction ranges considered here, such bound state formation is suppressed (see, $e.g.$, Ref. Gresham et al. (2018)). Accordingly, our heating signal arises from the energy deposition of individual dark matter particles. One may also worry that this long range force will introduce a new Jeans length for the Galactic dark matter halo. However, this does not occur for the range of couplings and mediator masses of the interaction we consider; see Appendix A for further discussion.
Finally, we note the broader motivation for considering dark matter self-interactions. These are well-motivated on both observational and theoretical grounds. For example, the wide diversity of inferred dark matter halo profiles on small scales is in tension with predictions from collisionless cold dark matter simulations (see $e.g.$ Ref. Tulin and Yu (2018) for a review). Self-interactions within the dark sector have been proposed as a resolution to this small-scale structure problem, while maintaining consistency with the successes of $\Lambda$CDM at large scales Kaplinghat et al. (2014); Elbert et al. (2015); Robertson et al. (2018). From a model-building perspective, long-range forces—mediated by either vector or scalar particles—remain viable. Scalar-mediated forces, in particular, are expected in scenarios such as those involving quintessence Farrar and Peebles (2004); Amendola and Tocchini-Valentini (2002). More recent studies have also explored collective effects of long-range forces in the dark sector Bogorad et al. (2023), though none to date have examined forces with the specific range considered in our work.
@@ II.2.1 Case I: Low Dark Matter Abundance, Low Strength Field
We first analyze the case of a single centrifugal barrier, when the dark potential $\Phi$ is sufficiently weak, and the force range $\lambda$ is sufficiently long, such that the $1/r$ factor in the dark potential $\Phi(r)$ limits capture. Because we are interested in a heating signal, we require that no centrifugal barrier be present at or beyond the surface of the celestial object, as otherwise this would prevent the dark matter from depositing its kinetic energy. This defines a maximum angular momentum, and therefore a maximum impact parameter $b_{\rm max}^{\rm dark}$, for which the point of closest approach matches the surface of the object ($i.e.$ Eq. (7) vanishes at the surface of the celestial object). Dark matter particles with angular momentum $b\leq b_{\rm max}^{\rm dark}$ thus contribute to heating, while those with $b>b_{\rm max}^{\rm dark}$ are unable to reach the object and deposit their energy. The full calculation is shown in Appendix C, however we predominantly work in the limit that $|\Phi|\ll m_{\chi}$. Under this approximation, we find that
@@ II.2.2 Case II: Sufficiently High Dark Matter Abundance, High Strength Field
As the number of dark matter particles inside the celestial body and therefore the dark potential grows over time, the maximum impact parameter also grows. While higher order terms in Eq. (11) need to be included in this regime, the fact that the maximum impact parameter $b_{\rm max}^{\rm dark}$ increases with increasing dark potential $\Phi$ can already be seen from Eq. (11). As the dark potential has a Yukawa scaling, its strength is exponentially suppressed at distances larger than the force range $\lambda$. Therefore, once a large amount of dark matter is accumulated and the impact parameter grows past the force range, there is eventually a shutoff where exponential growth no longer occurs, and the growth of the dark matter number becomes sub-linear. This effect is taken into account via a second centrifugal barrier, and is only relevant in the large dark matter number regime; we derive it in Appendix C. However, for most of our parameter space of interest, and most of the evolution time, Eq. (11) is sufficient and we focus our analytic calculations in the following sections to Case I.
@@ II.3 Dark Matter Energy Loss
In addition to increasing the effective capture radius, the long range force will also significantly boost in-falling dark matter particles. The boost factor as measured in the rest frame of the surface of the celestial object is given by (see Appendix D)
@@ II.4 Capture Probability
is the transition cross section, which marks the transition from the single to multi scatter regime, and $N_{\rm SM}$ is the number of target particles inside the celestial body of interest. Note that we have assumed that the object’s density profile corresponds to a uniform sphere. When requiring a dark matter particle to pass through a smaller radius, $\frac{1}{\delta}R$ as opposed to $R$, the capture rate only decreases by the same factor $\frac{1}{\delta}$, due to the scaling with the impact parameter $b_{\rm max}$ once the dark potential dominates. As such, the majority of the brown dwarf mass residing within $\frac{1}{\delta}R$ will have only small quantitative effects on the coupling strength $\alpha$ ($\alpha\xrightarrow{\sim}\delta\alpha$) required for observable heating. We therefore expect to capture the same qualitative effects that would result from a more detailed density profile, which in any case is not a steep profile for realistic brown dwarfs.
@@ II.5 Accumulation Rate
where $v_{p}$ is the peak of the dark matter velocity distribution (see Appendix E for details). We can now readily integrate Eq. (22) by writing
@@
where we have assumed the initial condition $N_{\chi}(0)=0$. Note that this equation fully relies on using Eq. (11) for $b_{\rm max}^{\rm dark}$, and as such, only applies to Case I of Section II.2. This is an exponential accumulation phase, given the form of Eq. (27). At very late times (or very large coupling $\alpha$ relative to force length $\lambda$), eventually the exponential growth of $N_{\chi}$ is shut off; see Appendix C.
In our considered parameter space, heating occurs before an accumulated dark matter number greater than about $10^{-6}N_{\rm SM}$. Once the dark matter density becomes comparable to that of the Standard Model density, other relevant changes to the object’s behavior could occur. Furthermore, once the dark matter number is high enough, the dark potential becomes comparable to the dark matter mass, and additional physical effects may become relevant that are outside the scope of this work (see Appendix D). For these reasons, we analyze only the scenario where $|\Phi|\lesssim m_{\chi}$. We find observable heating before this point, and note that for stronger couplings there will certainly be some observable signal, even if different to that from an increase in temperature pointed out here.
@@ III.1 Instantaneous Heating
with $T_{\rm SM}$ the temperature expectation from Standard Model sources in the absence of dark matter heating. We have assumed in Eq. (32) that all of the DM kinetic energy goes into an increase in the SM temperature of the object. While some small amount of the energy can go into increasing the radius of the object at the edge of our parameter space, this is at most a correction that deviates from the expectation under constant radius by only a factor of $\lesssim\sqrt{1.5}$ Macintosh (2023); Croon et al. (2024). For the bulk of our parameter space, it is even smaller and negligible. Assuming a constant radius is therefore a good approximation for our parameters of interest.
@@ IV Exoplanet and Brown Dwarf Kinetic Heating
Equipped with a framework for dark kinetic heating in the presence of long-range forces, we now apply our calculation to an example class of low-escape velocity objects: exoplanets and brown dwarfs. As elucidated in the introduction, these objects are advantageous for detection due to their large surface areas, leading to larger luminosities for fixed temperatures, but they also have an extensive upcoming search program with $e.g.$ JWST, Roman, and Rubin Telescopes. Tests far into the Galactic center also may allow potential probes into the largely unknown dark matter distribution in our Milky Way Galaxy. However, they have previously only been studied in the context of dark matter annihilation heating Leane and Smirnov (2021); Benito et al. (2024). We will show now that the discovery potential for this search is greatly enhanced when considering broader classes of particle physics models that produce the dark kinetic heating effect. We will consider other low-escape velocity objects in Sec. V that, depending on the model parameters, may also be of complementary interest.
@@ IV.1 Sensitivity to Coupling Strength and Force Range
Figure 4: Dark kinetic heating of a 10 Gyr old, 55 Jupiter mass brown dwarf with Jupiter’s radius, at local position as a function of the force range $\lambda$, with fixed coupling strength $\alpha=2.85\times 10^{-27}$. The gray dashed line “bg” indicates a background temperature of 750K, the blue line shows the dark kinetic heating, and the orange line displays the total brown dwarf temperature.
Figure 4 demonstrates dark kinetic heating of a brown dwarf as a function of force range $\lambda$. Here we have again taken $m_{\chi}=1$ GeV for simplicity. We choose a fixed value of the coupling $\alpha=2.85\times 10^{-27}$ as we vary the force range. This value is chosen as the maximum value of $\alpha$ such that heating shown is in a regime where energy deposition dominantly causes a temperature increase, rather than radius increase, of the object. We emphasize that this coupling $\alpha$ is chosen for demonstrative purposes, and any $\alpha\gtrsim 2.85\times 10^{-27}$ would have strong heating signals as well. This plot shares features similar to Fig. 3, where the left side asymptotes to the same value in the limit of negligible dark potential, and then an exponential increase in heating occurs as we increase the force range. Unlike the case of increasing the coupling $\alpha$, however, heating eventually saturates to a set value as we increase the force range large enough. This is because after a certain range ($\lambda\sim\mathcal{O}(10^{7})$ km for the example parameters we have chosen), the long range force is already behaving as essentially infinite relative to other parameters of the theory.
@@ IV.4 Constraints from WISE and JWST Data on a Super-Jupiter Exoplanet
Intriguingly, for dark matter parameters that would give the exact amount of heating required to sustain the 250 K temperature, one could also question whether the Super-Jupiter WISE 0885-0714 is actually significantly older and represents a positive dark matter signal, for alternate dark matter parameters. Indeed, it is important to note that the mass, temperature, and age of this Super-Jupiter are not measured independently, but rather inferred based on spectral data Leggett et al. (2017). This inference includes Standard Model-only assumptions, and therefore would change under the addition of dark kinetic heating. In any case, limits can however be drawn by taking conservative choices for any inferred Standard Model properties. A true positive signal of dark kinetic heating may look like a statistically significant number of objects with temperature and luminosity measurements that do not match standard cooling theory, nor standard expectations for any other object of the same age. This has been established in another work that focused on DM-heating of main-sequence stars John et al. (2024), where a new population is expected to appear that is observationally inconsistent with the SM. To further distinguish a brown dwarf heated by dark matter from similar objects, such as a younger red dwarf, one can also use lithium concentrations to independently determine the age of the object Basri (1998). In addition, microlensing measurements can provide an independent measure of brown dwarf masses in some cases Han et al. (2017). Finally, observation of objects which increase with temperature towards the galactic center, in a way that mirrors an expected DM distribution, would be a smoking gun signal of dark matter heating.
@@
Figure 8: Cross section required for 99% of dark matter particles that pass through a 55 Jupiter mass brown dwarf with the radius of Jupiter to be captured. An example fixed dark matter number of $N_{\chi}=5\times 10^{-12}N_{\rm SM}$ is used here, where $N_{\rm SM}$ is the number of Standard Model nucleons in the object. A fixed force range $\lambda$ equal to 100 times the brown dwarf’s radius is also chosen. The dashed blue line represents dark interaction strengths $\alpha=0$, displaying the required cross section in the absence of a long range force, and the solid lines represent different values of $\alpha$. The values of $N_{\chi}$ and $\alpha$ are arbitrarily chosen to demonstrate the effect on $\sigma_{\rm max}$ of a long range force, as $N_{\chi}$ is a function of time for any chosen value of $\alpha$.
@@ IV.5 Scattering Cross Section Sensitivity
Lastly, note in Figs. 8 and 9 that we have ignored any potential evaporation effects. Evaporation occurs when the dark matter is up-scattered by the celestial body matter, such that it can overcome the object’s escape velocity and leave the object, removing any potential dark matter signal. While long-range self-interactions will mitigate evaporation once some dark matter has been accumulated, the rapid evaporation of very light dark matter might negate the possibility of any build-up at all. However, there are a wide range of scenarios where evaporation does not generally occur, even at very light dark matter masses. For example, when the field $\phi$ also couples to nucleons evaporation is greatly mitigated Acevedo et al. (2024b), and can be suppressed by multiple orders of magnitude below usual contact interaction expectations. We expand on this point in Appendix F. We therefore simply note that some regions of Fig. 9 might require more specific ingredients than the minimal toy model adopted here.